OpenStudy (anonymous):

a compound consists of 22% nickel how much nickel will there be in 500g of compound

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

can u please answer

hero (hero):

\[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\] \[\frac{\text{amount of nickel}}{\text{total compound}} = \frac{\text{% nickel in compound}}{100}\] \[\frac{\text{x}}{\text{500}} = \frac{22}{100}\]

OpenStudy (anonymous):

so wahts the answer @Hero

hero (hero):

\(x = \) amount of nickel in the compound. Solve for x